OpenStudy (anonymous):

What are the foci of the ellipse given by the equation 4x2 + 9y2 – 32x + 18y + 37 = 0?

OpenStudy (imstuck):

You have to group all the like terms together and then I'm thinking you also have to compete the square, too.

OpenStudy (imstuck):

Do this:

OpenStudy (anonymous):

could you walk me through it please

OpenStudy (imstuck):

\[4x ^{2}-32x+9y ^{2}+18y=-37\]That's the beginnings of the proper set-up for this ellipse. Any ellipse, for that matter! Yes, I will walk you through it. See what I have done thus far?

OpenStudy (imstuck):

Now group the x terms in parenthesis and then the y terms in parenthesis. Like this (getting ready to complete the square).

OpenStudy (imstuck):

\[(4x ^{2}-32x)+(9y ^{2}+18y)=-37\]In order to complete the square you have to have a coefficient of 1 in front of both the x^2 and the y^2 terms. So factor out what you need to in order to make the coefficients a 1:

OpenStudy (imstuck):

\[4(x ^{2}-8x)+9(y ^{2}+2y)=-37\]Now it's time to complete the square.

OpenStudy (imstuck):

Take half the x term, square it, then add it in. Like this: Half of the x term (8) is 4; square 4 and get 16; add 16 into the parenthesis like this:

OpenStudy (imstuck):

\[4(x ^{2}-8x+16)+9(y ^{2}+2y)=-37\]Now do the same with the y's:

OpenStudy (imstuck):

\[4(x ^{2}-8x+16)+9(y ^{2}+2y+1)=-37\]BUT...

OpenStudy (imstuck):

Since you added them on to the left side of the equation, you also have to add them onto the right side. This is tricky, though, because you have a 4 and a 9 on the outside of the parenthesis. That means that when you add the 16 to the right side, when you pull it out of the parenthesis, you also have to keep in mind that you are multiplying by the 4. So it would look like this: (this is just the x terms, ok? I'm going to leave out the y's for just a sec. One crazy thing at a time, is what I say!)

OpenStudy (imstuck):

\[4(x ^{2}-8x+16)+...=-37+(4 * 16)\]I really hope that makes sense, cuz that's just the way it is!

OpenStudy (imstuck):

Now do it with the y's as well. This is only the y's; then we will put it all together. I just don't think it will all fit into the equation editor; it's too long all at one time!

OpenStudy (imstuck):

\[...+9(y ^{2}+2y+1)=-37+(4*16)+(9*1)\]

OpenStudy (imstuck):

Now you can reduce those parenthesis, because they are now perfect square binomials:

OpenStudy (imstuck):

\[4(x-4)^{2}+9(y+1)^{2}=-37 +64+9\]That right side equals\[4(x-4)^{2}+9(y+1)^{2}=36\]

OpenStudy (imstuck):

We all know that an ellipse has to = 1, so divide the equation by 36 to get a 1 on the right side. That looks like this:

OpenStudy (imstuck):

\[\frac{ 4(x-4)^{2} }{ 36 }+\frac{ 9(y+1)^{2} }{ 36 }=1\]Reduce between the 4 & 36 and the 9 & 36 to get this:

OpenStudy (imstuck):

\[\frac{ (x-4)^{2} }{ 9 }+\frac{ (y+1)^{2} }{ 4 }=1\]That is your ellipse now. The rule with ellipses is that a is always greater than b, and whichever variable the a is under --the x or the y--is the major axis. Here the major axis is the x axis. a^2 = 9 and a = 3; b^2 = 4 and b = 2.

OpenStudy (imstuck):

The equation for the foci is a^2 - b^2 = c^2. Or 9 - 4 = c^2 or c = sq rt of 5. Because this is an ellipse with the major vertices on the x axis, the foci are also on the x axis, which puts your sq rt 5 in the x position of your coordinates:

OpenStudy (imstuck):

\[(+\sqrt{5},0)(-\sqrt{5},0)\]Those are the foci!!!! Hope you understood that ok. I'm sure it's not the first time you saw this! Take it slow and read each step completely. It's a long and tedious process, but it is the ONLY way to get the job done! Any other questions, just tag me or someone else who loves doing these as much as I do, and we'll get you taken care of!

OpenStudy (anonymous):

(4, –1 ±square root of 5) (4, –1 ±square root of 13) (4 ±square root of 5 , –1) (4 ±square root of 5 , 1) Are the answer choices

OpenStudy (imstuck):

Let me get back to you o that in just a bit.

OpenStudy (imstuck):

I forgot to take into consideration the center is not at the origin. The center is at (4, -1). So that means take your foci of \[(\sqrt{5},0)\]and add the center's coordinates into it like this:

OpenStudy (imstuck):

\[(\sqrt{5}+4,0-1)=(4+\sqrt{5},-1)\]and the other one, the one with the negative sq rt 5 is this:

OpenStudy (imstuck):

\[(4-\sqrt{5},-1)\]THOSE are your foci!!!

OpenStudy (imstuck):

That is choice c (or the third one down).

OpenStudy (anonymous):

ok thank you, do you mind helping me with two more quaestions

OpenStudy (imstuck):

ok

OpenStudy (imstuck):

I can help you!